Question 1

The probabilities of events A and B are 0.5 and 0.1, respectively.  The probability that both A and B occur is 0.20. What is the probability of either A or B occuring?

a.

0.40

b.

0.1

c.

0.50

d.

0.60
Answer #1

The probabilities of events A is 0.5

The probabilities of events Bis 0.1

The probability that both A and B occur is 0.20.

P(AorB)=P(A)+P(B)-P(A andB)=0.5+0.1-0.2=0.4.

answer is a.0.4